2541234:

Isabel is thinking of two numbers. Their greatest common factor is 14. Their least common multiple is 168. One of the number is 56. what is the other number. A. 84 B. 42 c. 28 d. 8 e. 7

surjithayer:

second number\[=\frac{ 14 \times 168 }{ 56 }=14\times 3=42\] second number=42